Solve the differential equation (x +4) dy/dx = x(y-3), given that y=4 when x= -3.


1
Expert's answer
2022-05-02T13:32:09-0400

Solution

Given equation is separable differential equation. So

dy/(y-3) = xdx/(x+4)

dy/(y-3) = dx – 4dx/(x+4)

ln|y-3| = x – 4ln|x+4| + C

y-3 = exeC/(x+4)4  

y = 3 + exeC/(x+4)4  

If y=4 when x= -3 then 4 = 3 + eC e-3 => eC = e3  => y(x) = 3 + ex+3/(x+4)4    

Answer

y(x) = 3 + ex+3/(x+4)4